

********************************************************************
UARS WEEKLY STATUS REPORT - 02 APR 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM gathered its normal solar, offset, and occultation data without
apparent problems. The data from the new working channel continues to
indicate that it is working as expected.
Reprocessing of the SUSIM data to version 20 continues.
***********
PEM
***********
PEM/AXIS was activated March 30. While AXIS is activated, measured
X-rays are inverted to determine the incident electron population and
this electron population is used to determine the amount of energy
deposited in the atmosphere by electrons and bremsstrahlung X-rays.
PEM is active under PMON control. Energy processor #6 for the
HEPS-1 unit has been found to not activate correctly. Data affected
are from the upper energy EE channel for Head-2 of HEPS-1 (that is the
+15 deg electrons above 300 keV). Tests are being conducted on this
energy processor and other energy processors to determine conflicts.
Measurements from this sensor will be activated as soon as possible.
The PEM team and FOT are convinced that the entanglement of energy
processor #6 is related to OBC control and not a failure of that
energy processor. Early attempts at adjusting OBC command timing
have not been successful. The FOT and the PEM team is attempting to
refine their diagnostics and confine the problem.
While under PMON control, the PEM instrument computes energy deposition
from the particle spectrometers.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
The instrument is now off as scheduled. MLS is scheduled to begin north viewing
operations on Apr 4 and operate though Apr 10.
***********
HRDI
***********
HRDI is currently operating with a new set of tuning coefficients,
which have successfully increased the signal levels HRDI detects.
Unfortunately, these new coefficients will require us to determine new
line-of-sight zero wind positions and sensitivity constants, which we will need
to delivery to the CDHF before the HRDI level 3a results after uars day 2697
reach the same accuracy as previous data.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
2 days CLA (7.0) ENG (L1-L3) (8R-8C=0L)
50 days PEM (3.32) PROD (OA-L2) (760R-378C-1F=381L)
12 days SUS (7.0) PROD (677R-677C=0L)
11 days SUS (7.1) PROD (L3) (11R-11C=0L)
193 days WIN (5.11) PROD (LOG) (2014R-416C=1598L)
61 days WIN (5.11) PROD (L1-L3) (2014R-74C-1F=1939L)
Preventive Maintenance on the VAX cluster and peripherals is tentatively
scheduled for 04/06/99.
The CDHF is scheduled to participate in an End to End Y2K test
in mid-April. This test will require a system clock time shift and an
interruption to normal operations. A meeting is being scheduled to
further discuss the safest method of completing this task with the
least impact to UARS services. The CDHF is also investigating
a date similation product.
DAAC data created for January 1999 has been loaded into the DAAC transfer
tables for HALOE, NMC, PEM, and UKMO.
CDHF pending software upgrades: DEC AXP Fortran 6.5, DEC AXP OPEN
VMS 6.2, OpenIngres
SOLSTICE Wallclock Time Investigation.
VAX and ALPHA/debug version of the SOLSTICE step 2 program ran successfully
on the development system. Plan - build an operational version using Debug
units and run on the operational system Alpha next week.
***********
FOT
***********
REPORTING PERIOD
This report covers 20 March 1999 (Orbit 41095, GMT Day 079 through 26 March
1999.
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 083/2301 41169 Off
The beta angle ranged between the angles of 33.4 degrees to 7.1 degrees for
this report period. The beta angle is now decreasing toward a minimum
angle of 0 degrees on 28 March 1999 (DOY 087).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.9 to 1.2 degrees C. Additional battery performance data
for each battery on 19 March 1999 and 26 March 1999 are provided in
Appendix D.
The clock error ranged from -7.0 to 9.0 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS 00
% OUT-OF-LIMITS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 00
% HALOE 01
% PEM 02
% SUSIM 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 01
Two (2) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 99/054 - 2287.496948
Transponder B: 99/055 - 2287.498731
GROUND SYSTEM OPERATION
No (0) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 99/084 show the observatory data loss to be
22 hours, 46 minutes, 35.700 seconds (an increase of 0.672 seconds since
last report period). This is a 0.0346 percent data loss which equals a
99.9654 percent data capture for the mission.
7 The increase of 0.672 seconds occurred on Day 079 and was due to a
data hit.
The UARS Test and Training Simulator (UTTS) has been repaired and was
declared operational on 18 March 1999. The FOT conducted testing of the
UTTS through MSOCC. A few line configuration issues were identified by the
COMM MGR and corrected. We believe these to have been procedural problems
not related to any UTTS hardware problems.
Two more instances of three command-word corruption of the Ground Reference
Image occurred on 28 March 1999. This time the incorrectly modified words
were in the ATC location of memory. Analysis of the dump compare data
demonstrated that the miscompares were outside of the area just loaded.
While once again there was no operational impact from this recurring
problem, it does undermine the FOT's confidence in the ground reference
image as a diagnostic tool.
The FOT is still awaiting the delivery of the two remaining Y2K test
planning aids from FDF. These planning aids were originally requested by
SOLSTICE Operations via the Mission Planning Group.
The FOT experienced several failed attempts at transferring the daily CMS
generated loads to DFLX this week. Instead of each load in the
POCCATLOG.DAT file transferring successfully, one or more of the loads in
each daily group would result in a fatal error. Investigation by CMF found
that there is a faulty disk on the DFLX system, which resulted in the
occasional unsuccessful load transfer. On the days in which the load
transfer process resulted in an error, the FOT was required to re-transmit
the loads one or two more times (until the transfer process was
successful). CMF is currently working on a plan to replace the faulty disk
on the DFLX system, as well as a plan for getting the backup DFLX system
(DFLX2) into an operational status.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT continues to support the PEM team with diagnosis of the Energy
Processor 6 status error. PEM EP6 testing requires a new RTS 31 and
possibly other minor modifications for the next step in testing. It has
been determined that whenever PDP-10 is enabled and activates, EP6 becomes
disabled. The command sequences in PDP-10 are correct. The next test will
move those three commands into an RTS.
The FOT continues to investigate the feasibility of a new PMON action to
change battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
YEAR 2000 OPERABILITY SUPPORT - No change
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
Planning continues for the simulator testing.
MEETINGS
FOT supported an IMOC meeting on 24 March 1999. The Operations Concept
production schedule was discussed and roughed in with labor estimates. The
team made assignments of issue papers related to various IMOC scenarios.
The FOT will participate in preparing several of these documents, leading
the effort on two of them.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver (forward-to-reverse) is scheduled for 28
March 1999 (DOY 087).
PERSONNEL CHANGES
None.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 26 March 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
41097 079/0339 99-111 PEM PEM TEST
41139 081/2138 99-112 PEM PEM PBD Test
41156 083/0217 99-113 SUSIM SUSIM Update
41169 083/2301 99-114 HALOE Power OFF
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 26 March 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 26 March 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
-No clock adjusts for this report period
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 26 March 1999
19 Mar 1999 - GMT Day 078 Beta = 33.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -22.4/-44.8 -3.68/-4.94 +0.0/ +0.0 16.8 0.000 0.0
2 +56.0/ 0.0 +5.15/+3.87 +29.6/-14.4 25.9 1.025 20.6
3 +44.8/-22.4 +3.87/+2.60 +28.8/-14.4 25.9 1.019 20.1
26 Mar 1999 - GMT Day 085 Beta = 7.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -33.6/-50.4 -3.68/-4.94 +0.0/ +0.0 16.8 0.000 0.0
2 +44.8/-11.2 +4.83/+3.55 +30.4/-13.6 25.6 1.016 20.6
3 +89.6/0.0 +3.87/+2.60 +30.4/-14.4 25.4 1.014 20.4
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 26 March 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-014 FOT 41189 085/0548 N/A CRC Errors in PBK
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-011 FOT see list Mar 99 14170 Generic Late Acq
99-014 FOT 41189 085/0548 N/A CRC Errors in PBK
SEVNTFW99/13